# Tony Lindeberg

### v1

@@ 1 Introduction
To handle such variabilities in image structures caused by variations in the viewing conditions, the notion of multi-parameter scale spaces has been proposed regarding the families of receptive fields (Lindeberg 2011; 2021). Specifically, from the desirable property of provable covariance properties of the families of spatial and/or spatio-temporal receptive fields under the appropriate families of geometric image transformations, the shapes of the spatial or spatio-temporal receptive fields ought to be expanded over the degrees of freedom of the corresponding image transformations (Lindeberg 2025b). Notably, such a notion has been proposed as a tentative model for variabilities of simple cells in the primary cortex of higher mammals (Lindeberg 2025c).
If one in an idealized vision system is to take the view that a rich set of receptive field responses is to be computed for a rich variety of shape parameters of the receptive fields, one can raise the question if this could be done in a more efficient manner than computing each receptive field response separately. Similarly, one may ask how receptive field responses computed for different values of the filter parameters could be related. For example, for the pure spatial scale-space representation defined from convolution with either isotropic or anisotropic affine Gaussian kernels, the underlying spatial smoothing filters obey cascade smoothing properties over spatial scales, implying that any representation at a coarser spatial scale can be computed by applying a filter or a set of filters to the representations at any finer scale.
@@
The subject of this paper, is to generalize the above mentioned ideas for the specific generalized Gaussian derivative model for visual receptive fields, initially derived in Lindeberg (2011; 2013) and then refined in Lindeberg (2016; 2021; 2023b; 2025b; 2025c). Using the algebraic properties of the idealized models for visual receptive fields according to this framework, we will derive closed-form expressions for derivatives with respect to the filter parameters in this model as well as explicit macroscopic cascade relations between receptive field responses for different parameter settings, roughly corresponding to a a combination of Lie group structures with multi-parameter semi-group structures, comprising 3 effective parameters in the purely spatial case and 4 or 6 effective parameters in the joint spatio-temporal case, although for special reasons of separating the degree of freedom corresponding to purely spatial scaling transformations, overparameterized over 4 or 7 parameters when using affine Gaussian kernels for the spatial smoothing operations.
This will result in (i) sets of infinitesimal generators and (ii) sets of macroscopic cascade smoothing properties over subsets of the parameter spaces for which the evolution properties can only performed in the positive parameter directions. For other subsets in the parameter space, the corresponding relationships between receptive field responses for different parameter settings will, however, be bidirectional.
@@
Specifically, in relation to the recently proposed tentative model for variabilities of simple cells in the primary cortex of higher mammals in Lindeberg (2025c), the presented results provide a detailed theoretical foundation for how to realize the alternative conceptual model in that paper, where the receptive field responses corresponding to simple cells are not computed for all spatial and/or temporal scales, but instead because of the semi-group properties with the spatial and the temporal scale parameters only computed at the finest spatial and/or temporal scales at each image position and temporal moment. Thereby, as considered as a possible design option for an idealized vision system, the image representations at higher layers in the visual hierarchy would not make use of explicit receptive field responses computed for simple cells for each spatial and/or temporal scale, but could instead compute equivalent receptive field responses at higher levels in the visual hierarchy, from the responses of the explicitly computed responses of simple cells as only implemented at the finest spatial and/or temporal scales.
Additionally, if an idealized vision system would be based on computing receptive field responses as corresponding to the output from simple cells, the relationships between receptive field responses for different values of the filter parameters constitute a theoretical foundation for how such responses could be computed in a hierarchical network structure, where receptive field responses are computed in a way that constitutes an extension to the cascade smoothing properties over the spatial and/or the temporal scale parameters, to enable a computationally more efficient implementation.
@@ 1.1 Structure of this article
The presentation is organized as follows: After describing relations to previous work in Section 2, Section 3 starts by reviewing the generalized Gaussian derivative model for visual receptive fields, including its covariance properties under important classes of geometric image transformations in terms of (i) spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations. From the closedness properties of the receptive field responses under these classes of geometric image transformation, we specifically explain why the aim of being able to match receptive field responses computed under different viewing conditions leads to the notion of multi-parameter filter banks of receptive fields, which motivates this study of theoretically showing how receptive field responses computed for different values of the filter parameters can be related. This section also defines the notation and the terminology in the way as it will be used in the forthcoming more technical sections.
Section 4 then begins the new theoretical work by deriving infinitesimal relationships between receptive field responses for subclass of receptive field models that are based on spatial or spatio-temporal smoothing solely based on Gaussian kernels over space or joint space-time. Section 5 then continues by deriving macroscopic relations between filter responses computed for different sets of filter parameters, based on either purely spatial or joint spatio-temporal cascade smoothing properties of the corresponding receptive field representations, defined from spatial or spatio-temporal smoothing based on solely using Gaussian smoothing kernels over space or joint space-time.
@@ 2 Relations to previous work
Concerning theoretical modelling of visual receptive fields, the regular Gaussian derivative model was initially proposed by Koenderink and van Doorn (1984; 1987; 1992) used for modelling biological receptive fields by Young (1987) and his co-workers in Young et al. (2001; 2001). This regular Gaussian derivative model has also been used as a component in more developed models of biological vision by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Wallis and Georgeson (2009), Hansen and Neumann (2008), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
@@
The generalization into receptive field families based on multi-parameter scale spaces goes back to the formulation of affine Gaussian scale space representation in Lindeberg and Gårding (1997) and early formulations of discrete scale-space representations over either purely spatial or joint spatio-temporal domains in Lindeberg (1997; 2001). These notions were then extended into the generalized Gaussian derivative model for either purely spatial or joint spatio-temporal domains in Lindeberg (2011; 2013; 2021) and with additional extensions concerning provable covariance properties of receptive field families in Lindeberg (2023b; 2025b; 2025c).
@@
The variabilities in receptive field shapes genererated by the affine Gaussian scale space have been used for computing more accurate estimates of local surface orientation from monocular och binocular cues by Lindeberg and Gårding (1997) and Rodríguez et al. (2018), for computing affine invariant image features for image matching under wide baselines by Baumberg (2000), Mikolajczyk and Schmid (2004), Mikolajczyk et al. (2005), Tuytelaars and van Gool (2004), Lazebnik et al. (2005), Rothganger et al. (2006; 2007), Lia et al. (2013), Eichhardt and Chetverikov (2018) andi Dai et al. (2020), for performing affine invariant segmentation by Ballester and González (1998), for constructing affine covariant SIFT descriptors by Morel and Yu (2009), Yu and Morel (2009) and Sadek et al. (2012), for modelling receptive fields in biological vision by Lindeberg (2013; 2021), for affine invariant tracking by Giannarou et al. (2013), and for formulating affine covariant metrics by Fedorov et al. (2015). This rich variety of application domains clearly demonstrates how the use of variabilities in the shapes of the receptive fields as implied by covariant receptive field representations (here with regard to spatial affine transformations) can be used for improving the performance of visual operations.
For modelling spatial receptive fields in vision, also Gabor models have been used by Marcelja (1980), Jones and Palmer (1987a; 1987b), Ringach (2002; 2004), Serre et al. (2007) and De and Horwitz (2021). The applicability of Gabor filters to spatio-temporal image data has, however, not been extensively explored, why we restrict ourselves to the generalized Gaussian derivative model in this treatment.
@@
With the in this paper combined treatment of 4 main classes of geometric image transformations involving (i) spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, the underlying aims of this work bear close similarities to approaches in the area of geometric deep learning (Bronstein et al. 2021, Gerken et al. 2023), which also consider visual processing operations that are covariant under wider classes of geometric image transformations.
@@ 3.1 Spatial and spatio-temporal smoothing kernels
In the case of a non-causal temporal domain, i.e., for pre-recorded video, for which the future can be accessed, the temporal smoothing kernel $h(t;\;\tau)$ in the spatio-temporal smoothing kernel $T(x,t;\;s,\Sigma,\tau,v)$ can be chosen as Gaussian kernel
$h(t;\;\tau)=g(t;\;\tau)=\frac{1}{\sqrt{2\pi}\sqrt{\tau}}\,e^{-t^{2}/2\tau^{2}}.$ (3)
@@
for some initial temporal scale level $\tau_{0}\in{\mathbb{R}}_{+}$, where the distribution parameter $c$ is for common purposes of implementation often chosen as $c=\sqrt{2}$ or $c=2$.
In fact, it is shown in an axiomatic way in Lindeberg (2011), that the forms of the spatial smoothing kernel (1) and the spatio-temporal smoothing kernel (2) are uniquely determined, given natural symmetry requirements on a visual front end, with conceptual extensions of those ideas to a time-causal temporal domain in Lindeberg (2016; 2021). The underlying symmetry requirements used for the axiomatic derivations reflect symmetry properties of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over multiple spatial and temporal scales, see Lindeberg (2011; 2021) for further details.
@@
Figure 2: Non-causal joint spatio-temporal receptive fields over a 1+1D spatio-temporal domain in terms of the mixed first-order spatial derivative and the first-order velocity-adapted temporal derivative of the form $T_{x\bar{t}}(x,t;\;s,\tau,v)$ according to (23) of the product of a velocity-adapted 1-D Gaussian kernel over the spatial domain and the non-causal temporal kernel over the temporal domain according to (3). The spatio-temporal receptive fields are shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ and the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ in dimensions of $\mbox{length}]$ and $\mbox{time}]$. (Horizontal axes: Spatial image coordinate $x\in[-10,10]$. Vertical axes: Temporal variable $t\in[-4,4]$.)
@@
Figure 3: Time-causal joint spatio-temporal receptive fields over a 1+1D spatio-temporal domain in terms of the mixed first-order spatial derivative and the first-order velocity-adapted temporal derivative of the form $T_{x\bar{t}}(x,t;\;s,\tau,v)$ according to (23) of the product of a velocity-adapted 1-D Gaussian kernel over the spatial domain and the time-causal limit kernel over the temporal domain according to (4). The spatio-temporal receptive fields are shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ and the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ in dimensions of $\mbox{length}]$ and $\mbox{time}]$. (Horizontal axes: Spatial image coordinate $x\in[-10,10]$. Vertical axes: Temporal variable $t\in[0,8]$.)
@@ 3.2 Receptive field families in terms of spatial and spatio-temporal derivatives
For defining either purely spatial or joint spatio-temporal receptive fields to be used for processing visual image data, the either purely spatial or joint spatio-temporal smoothing operations are, in turn, combined with either purely spatial differential operators ${\cal D}_{x}$ or joint spatio-temporal differentiation operators ${\cal D}_{x,t}$, to give rise to spatial receptive fields applied to spatial image data $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ of the form
${\cal R}\,f(\cdot)={\cal D}_{x}(T(\cdot;\;s,\Sigma)*f(\cdot))$ (7)
@@
Regarding the choice of what spatial differentiation operators to use, common choices include directional derivative operators in the orientation $\varphi\in[-\pi,\pi]$ of the form
@@
Concerning joint spatio-temporal derivative operators, a natural choice is to combine the above purely spatial derivative operators ${\cal D}_{x}$ with velocity-adapted temporal derivative operators $\partial_{\bar{t}}^{n}$ according to
@@
${\cal D}_{x,t}={\cal D}_{x}\,\partial_{\bar{t}}^{n}.$ (13)

### v2

@@ 1 Introduction
To handle such variabilities in image structures, caused by variations in the viewing conditions, the notion of multi-parameter scale spaces has been proposed regarding the families of receptive fields (Lindeberg 2011; 2021). Specifically, from the desirable property of provable covariance properties of the families of spatial and/or spatio-temporal receptive fields under the appropriate families of geometric image transformations, the shapes of the spatial or spatio-temporal receptive fields ought to be expanded over the degrees of freedom of the corresponding image transformations (Lindeberg 2025c). Notably, such a notion has been proposed as a tentative model for variabilities of simple cells in the primary cortex of higher mammals (Lindeberg 2025d).
If one in an idealized vision system is to take the view that a rich set of receptive field responses is to be computed for a rich variety of shape parameters of the receptive fields, one can raise the question if this could be done in a more efficient manner than computing each receptive field response separately. Similarly, one may ask how receptive field responses computed for different values of the filter parameters could be related. For example, for the pure spatial scale-space representation, defined from convolution with either isotropic or anisotropic affine Gaussian kernels, the underlying spatial smoothing filters obey cascade smoothing properties over spatial scales, implying that any representation at a coarser spatial scale can be computed by applying a filter or a set of filters to the representations at any finer scale.
@@
The subject of this paper, is to generalize the above mentioned ideas for the specific generalized Gaussian derivative model for visual receptive fields, initially derived in Lindeberg (2011; 2013a) and then refined in Lindeberg (2016; 2021; 2023b; 2025c; 2025d). Using the algebraic properties of the idealized models for visual receptive fields according to this framework, we will derive closed-form expressions for derivatives with respect to the filter parameters in this model as well as explicit macroscopic cascade relations between receptive field responses for different parameter settings, roughly corresponding to a combination of Lie group structures with multi-parameter semi-group structures, comprising 3 effective parameters in the purely spatial case and 4 or 6 effective parameters in the joint spatio-temporal case, although for special reasons of separating the degree of freedom corresponding to purely spatial scaling transformations, overparameterized over 4 or 7 parameters when using affine Gaussian kernels for the spatial smoothing operations.
This will result in (i) sets of infinitesimal generators and (ii) sets of macroscopic cascade smoothing properties over subsets of the parameter spaces, for which the evolution properties can only performed in the positive parameter directions. For other subsets in the parameter space, the corresponding relationships between receptive field responses for different parameter settings will, however, be bidirectional.
@@
Specifically, in relation to the recently proposed tentative model for variabilities of simple cells in the primary cortex of higher mammals in Lindeberg (2025d), the presented results provide a detailed theoretical foundation for how to realize the alternative conceptual model in that paper, where the receptive field responses corresponding to simple cells are not computed for all spatial and/or temporal scales, but instead because of the semi-group properties with the spatial and the temporal scale parameters only computed at the finest spatial and/or temporal scales at each image position and temporal moment. Thereby, as considered as a possible design option for an idealized vision system, the image representations at higher layers in the visual hierarchy would not make use of explicit receptive field responses computed for simple cells for each spatial and/or temporal scale, but could instead compute equivalent receptive field responses at higher levels in the visual hierarchy, from the responses of the explicitly computed responses of simple cells as only implemented at the finest spatial and/or temporal scales.
Additionally, if an idealized vision system would be based on computing receptive field responses as corresponding to the output from simple cells, then the relationships between receptive field responses for different values of the filter parameters constitute a theoretical foundation for how such responses could be computed in a hierarchical network structure, where receptive field responses are computed in a way that constitutes an extension to the cascade smoothing properties over the spatial and/or the temporal scale parameters, to enable a computationally more efficient implementation.
@@ 1.1 Structure of this article
The presentation is organized as follows: After describing relations to previous work in Section 2, Section 3 starts by reviewing the generalized Gaussian derivative model for visual receptive fields, including its covariance properties under important classes of geometric image transformations in terms of (i) spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations. From the closedness properties of the receptive field responses under these classes of geometric image transformation, we specifically explain why the aim of being able to match the receptive field responses computed under different viewing conditions leads to the notion of multi-parameter filter banks of receptive fields, which motivates this study of theoretically showing how receptive field responses computed for different values of the filter parameters can be related. This section also defines the notation and the terminology in the way as it will be used in the forthcoming more technical sections.
Section 4 then begins the new theoretical work by deriving infinitesimal relationships between receptive field responses for the subclass of receptive field models, that are based on spatial or spatio-temporal smoothing solely based on Gaussian kernels over image space or joint space-time. Section 5 then continues by deriving macroscopic relations between filter responses computed for different sets of filter parameters, based on either purely spatial or joint spatio-temporal cascade smoothing properties of the corresponding receptive field representations, defined from spatial or spatio-temporal smoothing based on solely using Gaussian smoothing kernels over image space or joint space-time.
@@ 2 Relations to previous work
Concerning theoretical modelling of visual receptive fields, the regular Gaussian derivative model was initially proposed by Koenderink and van Doorn (1984; 1987; 1992) and used for modelling biological receptive fields by Young (1987) and his co-workers in Young et al. (2001; 2001). This regular Gaussian derivative model has also been used as a component in more developed models of biological vision by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Wallis and Georgeson (2009), Hansen and Neumann (2008), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
@@
The generalization into receptive field families based on multi-parameter scale spaces goes back to the formulation of affine Gaussian scale space representation in Lindeberg and Gårding (1997) and early formulations of discrete scale-space representations over either purely spatial or joint spatio-temporal domains in Lindeberg (1997; 2001). These notions were then extended into the generalized Gaussian derivative model for either purely spatial or joint spatio-temporal domains in Lindeberg (2011; 2013a; 2021) and with additional extensions concerning provable covariance properties of receptive field families in Lindeberg (2023b; 2025c; 2025d).
@@
The variabilities in receptive field shapes genererated by the affine Gaussian scale space have been used for computing more accurate estimates of local surface orientation from monocular and binocular cues by Lindeberg and Gårding (1997) and Rodríguez et al. (2018), for computing affine invariant image features for image matching under wide baselines by Baumberg (2000), Mikolajczyk and Schmid (2004), Mikolajczyk et al. (2005), Tuytelaars and van Gool (2004), Lazebnik et al. (2005), Rothganger et al. (2006; 2007), Lia et al. (2013), Eichhardt and Chetverikov (2018) and Dai et al. (2020), for performing affine invariant segmentation by Ballester and González (1998), for constructing affine covariant SIFT descriptors by Morel and Yu (2009), Yu and Morel (2009) and Sadek et al. (2012), for modelling receptive fields in biological vision by Lindeberg (2013a; 2021), for affine invariant tracking by Giannarou et al. (2013), and for formulating affine covariant metrics by Fedorov et al. (2015). This rich variety of application domains clearly demonstrates how the use of variabilities in the shapes of the receptive fields, as implied by covariant receptive field representations (here with regard to spatial affine transformations), can be used for improving the performance of visual operations.
For modelling spatial receptive fields in vision, also Gabor models have been used by Marcelja (1980), Jones and Palmer (1987a; 1987b), Ringach (2002; 2004), Serre et al. (2007), Baspinar et al. (2012; 2018; 2020), De and Horwitz (2021) and others. The use of Lie symmetries for deriving receptive field shapes based on Gabor filters under the SE(2) group has been studied by Citti and Sarti et al. (2006) and under the 2-D affine group by Sarti et al. (2008). Modelling of spatio-temporal receptive fields using Gabor functions has been proposed by Cocci et al. (2011) and Barbieri et al. (2014).
Both the complex-valued Gabor functions used in the Gabor model and the real-valued Gaussian functions underlying the Gaussian derivative model minimize the uncertainty relation (Barbieri et al. 2012, Lindeberg 2013b).
@@
With the combined treatment of the 4 main classes of geometric image transformations considered in this paper, involving (i) spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, the underlying aims of this work bear close similarities to approaches in the area of geometric deep learning (Bronstein et al. 2021, Gerken et al. 2023), which also consider visual processing operations that are covariant under wider classes of geometric image transformations.
@@ 3.1 Spatial and spatio-temporal smoothing kernels
In the case of a non-causal temporal domain, i.e., for pre-recorded video, for which the future can be accessed, the temporal smoothing kernel $h(t;\;\tau)$ in the spatio-temporal smoothing kernel $T(x,t;\;s,\Sigma,\tau,v)$ can be chosen as a Gaussian kernel
$h(t;\;\tau)=g(t;\;\tau)=\frac{1}{\sqrt{2\pi}\sqrt{\tau}}\,e^{-t^{2}/2\tau}.$ (3)
@@
for some initial temporal scale level $\tau_{0}\in{\mathbb{R}}_{+}$. For common purposes of implementation, the distribution parameter $c$ is often chosen as $c=\sqrt{2}$ or $c=2$.
In fact, it is shown in an axiomatic way in Lindeberg (2011), that the forms of the spatial smoothing kernel (1) and the spatio-temporal smoothing kernel (2) are uniquely determined, given natural symmetry requirements on a visual front end, with conceptual extensions of those ideas to a time-causal temporal domain in Lindeberg (2016; 2021). The underlying symmetry requirements used for the axiomatic derivations are based on symmetry properties of the environment, in combination with internal consistency requirements, to guarantee theoretically well-founded treatment of image structures over multiple spatial and temporal scales, see Lindeberg (2011; 2021) for further details.
@@
Figure 2: Non-causal joint spatio-temporal receptive fields over a 1+1D spatio-temporal domain in terms of the mixed first-order spatial derivative and the first-order velocity-adapted temporal derivative of the form $T_{x\bar{t}}(x,t;\;s,\tau,v)$ according to (24) as the product of a velocity-adapted 1-D Gaussian kernel $g_{1D}(x-v\,t;\;s)$ over the spatial domain and the non-causal Gaussian kernel $h(t;\;\tau)=g(t;\;\tau)$ over the temporal domain according to (3). The spatio-temporal receptive fields are shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ and the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ in dimensions of $\mbox{length}]$ and $\mbox{time}]$. (Horizontal axes: Spatial image coordinate $x\in[-10,10]$. Vertical axes: Temporal variable $t\in[-4,4]$.)
@@
Figure 3: Time-causal joint spatio-temporal receptive fields over a 1+1D spatio-temporal domain in terms of the mixed first-order spatial derivative and the first-order velocity-adapted temporal derivative of the form $T_{x\bar{t}}(x,t;\;s,\tau,v)$ according to (24) as the product of a velocity-adapted 1-D Gaussian kernel $g_{1D}(x-v\,t;\;s)$ over the spatial domain and the time-causal limit kernel $h(t;\;\tau)=\psi(t;\;\tau,c)$ over the temporal domain according to (4) with the distribution parameter set to $c=2$. The spatio-temporal receptive fields are shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ and the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ in dimensions of $\mbox{length}]$ and $\mbox{time}]$. (Horizontal axes: Spatial image coordinate $x\in[-10,10]$. Vertical axes: Temporal variable $t\in[0,8]$.)
@@ 3.2 Receptive field families in terms of spatial and spatio-temporal derivatives
For defining either purely spatial or joint spatio-temporal receptive fields, to be used for processing visual image data, the either purely spatial or joint spatio-temporal smoothing operations are, in turn, combined with either purely spatial differential operators ${\cal D}_{x}$ or joint spatio-temporal differentiation operators ${\cal D}_{x,t}$, to give rise to spatial receptive fields applied to spatial image data $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ of the form
${\cal R}\,f(\cdot)={\cal D}_{x}(T(\cdot;\;s,\Sigma)*f(\cdot)),$ (7)
@@
Regarding the choice of what spatial differentiation operators ${\cal D}_{x}$ to use, common choices include directional derivative operators in the orientation $\varphi\in[-\pi,\pi]$ of the form
@@
Concerning joint spatio-temporal derivative operators ${\cal D}_{x,t}$, a natural choice is to combine the above purely spatial derivative operators ${\cal D}_{x}$ with velocity-adapted temporal derivative operators $\partial_{\bar{t}}^{n}$ according to
@@
${\cal D}_{x,t}={\cal D}_{x}\,\partial_{\bar{t}}^{n},$ (13)
or alternatively using space-time separable (not velocity-adapted) spatio-temporal derivative operators of the form
${\cal D}_{x,t}={\cal D}_{x}\,\partial_{t}^{n}.$ (14)